OpenStudy (anonymous):

Find an explicit rule for the nth term of a geometric sequence where the second and fifth terms are -12 and 768, respectively. an = 3 • (-4)n + 1 an = 3 • 4n - 1 an = 3 • (-4)n - 1 an = 3 • 4n

OpenStudy (anonymous):

@RadEn

OpenStudy (campbell_st):

ok... so the basic formula for a term in a geometric sequence is \[a_{n} = ar^{n -1}\] here is what you know 2nd term \[-12 = a r^{2 -1} ...or...... -12 = ar\] the 5th term \[768 = a r^{5 -1} ....or...... 768 = ar^4\] this can be written using the 2nd term \[768 = ar \times r^3\] substitute the value of the 2nd term \[768 = -12 \times r^3\] you can now solve for r.... when you get r, substitute it into the 2nd term to find a. then just substitute r and a into the general formula above to get your answer. hope it helps